# Anytime-valid, Bayes-assisted, Prediction-Powered Inference

### v1

@@ Abstract
Given a large pool of unlabelled data and a smaller amount of labels, prediction-powered inference (PPI) leverages machine learning predictions to increase the statistical efficiency of standard confidence interval procedures based solely on labelled data, while preserving their fixed-time validity. In this paper, we extend the PPI framework to the sequential setting, where labelled and unlabelled datasets grow over time. Exploiting Ville’s inequality and the method of mixtures, we propose prediction-powered confidence sequence procedures that are valid uniformly over time and naturally accommodate prior knowledge on the quality of the predictions to further boost efficiency. We carefully illustrate the design choices behind our method and demonstrate its effectiveness in real and synthetic examples.
@@ 1 Introduction
Increasing the sample size of an experiment is arguably the single simplest way to improve the precision of the statistical conclusions drawn from it. However, in many fields – such as healthcare, finance, and social sciences – obtaining labelled data is often costly and time-consuming. In these settings, using machine learning models to impute additional labels represents a tempting alternative to expensive data collection, albeit at the risk of introducing bias. Prediction-powered inference (PPI) (Angelopoulos et al., 2023a) is a recently introduced framework for valid statistical inference in the presence of a small labelled dataset and a large number of unlabelled examples paired with predictions from a black-box model.
@@
$\theta^{\star}=\underset{\theta\in\mathbb{R}}{\arg\min}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{E}[\ell_{\theta}(X,Y)],$ (1)
where $\ell_{\theta}(x,y)$ is a convex loss function parameterised by $\theta\in\mathbb{R}$. As an example, the mean $\theta^{\star}=\mathbb{E}[Y]$ is the estimand induced by the squared loss $\ell_{\theta}(x,y)=(\theta-y)^{2}/2$. For $t=1,2,\ldots$, we observe a sequence of independent random variables $Z_{t}$, either drawn from $\mathbb{P}$ (labelled sample) or from $\mathbb{P}_{X}$ (unlabelled sample), and we are provided with a black-box prediction rule $f$, which maps any input $x$ to a prediction $f(x)$. Let $(X_{i},Y_{i})_{i\geq 1}$ and $(\widetilde{X}_{j})_{j\geq 1}$ denote the subsequence of labelled and unlabelled samples, respectively. For $n=1,2,\ldots$, let $N_{n}$ denote the number of unlabelled samples observed before the $n$th labelled one, and assume that $N_{n}\geq n$, with $N_{n}\gg n$ in typical settings. PPI constructs an (asymptotic) $1-\alpha$ confidence interval (CI) $\mathcal{C}_{\alpha,n}^{\mathrm{pp}}$ for $\theta^{\star}$, that exploits the auxiliary information encoded in $f$. To this end, under mild assumptions, $\theta^{\star}$ can be expressed as the solution to
@@
where $\ell_{\theta}^{\prime}$ is a subgradient of $\ell_{\theta}$ with respect to $\theta$. The quantity $g_{\theta}$ in Equation 2 can be decomposed into the sum $g_{\theta}=m_{\theta}+\Delta_{\theta}$ of two terms,
@@
where $m_{\theta}$ represents a measure of fit of the predictor, while $\Delta_{\theta}$, called the rectifier, accounts for the discrepancy between the predicted outputs $f(X)$ and the true labels $Y$. If $\mathcal{C}^{g}_{\alpha,\theta,n}$ is a $(1-\alpha)$ confidence interval for $g_{\theta}$, then the PPI confidence interval $\mathcal{C}_{\alpha,n}^{\mathrm{pp}}$, defined as
@@
also achieves the desired coverage, i.e. $\Pr(\theta^{\star}\in\mathcal{C}_{\alpha,n}^{\mathrm{pp}})\geq 1-\alpha$. The construction of $\mathcal{C}^{g}_{\alpha,\theta,n}$ naturally relies on estimating $g_{\theta}$, for which PPI defines an estimator that leverages both the unlabelled data and the prediction rule $f$. The resulting method outperforms standard CI procedures based on the labelled data alone when $f$ is sufficiently accurate and $N_{n}\gg n$. Intuitively, this is due to the fact that, in this case, $\Delta_{\theta}$ is close to zero, while $m_{\theta}$ can be estimated with low variance from the unlabelled data.
Crucially, coverage of the PPI CI (4) is guaranteed only at a fixed time, i.e. for a labelled sample size $n$ fixed in advance. This is undesirable in many practical settings – such as online learning, real-time monitoring, or sequential decision-making – where it is essential to continuously draw conclusions as new data arrives. In this work, we address this by proposing an anytime-valid extension of the PPI CI (4). That is, we define a confidence sequence $(\mathcal{C}_{\alpha,n}^{\mathrm{avpp}})_{n\geq 1}$, satisfying the stronger coverage guarantee
@@
while still taking advantage of the prediction rule $f$. In doing so, similarly to standard PPI, we construct a confidence sequence $(\mathcal{C}^{g}_{\alpha,\theta,n})_{n\geq 1}$ for $g_{\theta}$ and define $\mathcal{C}_{\alpha,n}^{\mathrm{avpp}}$ through Equation 4
for $n\geq 1$. While our approach is agnostic to the specific form of the confidence sequence $(\mathcal{C}^{g}_{\alpha,\theta,n})_{n\geq 1}$, we mainly focus on asymptotic confidence sequences (Waudby-Smith et al., 2024b), as they represent a versatile time-uniform equivalent to standard CLT-based CIs that apply to the PPI framework above in full generality. Moreover, being based on the method of mixtures (Ville, 1939; Robbins, 1970; Lai, 1976), they can easily accommodate prior information on the quality of the prediction model $f$. In particular, by means of a zero-centered prior on the rectifier $\Delta_{\theta}$, we can obtain tighter confidence sequences when the predictions are good, extending the fixed-time Bayes-assisted approach of Cortinovis and Caron (2025).
The remainder of the paper is organised as follows. Section 2 reviews related work. Section 3 provides background on (asymptotic) confidence sequences and discusses how prior information may be easily incorporated into their construction. Section 4 presents PPI in the context of control-variate estimators, whose asymptotic properties are crucial for our approach to anytime-valid, Bayes-assisted, PPI, which is described in Section 5. Section 6 demonstrates the benefits of our method on synthetic and real data. Finally, Section 7 discusses limitations and further extensions of our approach. Proofs and additional experiments are provided in the Supplementary Material.
@@ 2 Related Work
Prediction-powered inference (PPI) was introduced by Angelopoulos et al. (2023a) as a general framework for valid statistical inference with black-box machine-learning predictors, and was later extended in (Angelopoulos et al., 2023b). Closely related ideas appear in the literatures on semi-supervised inference, missing-data methods, survey sampling, and double machine learning (Robins and Rotnitzky, 1995; Särndal et al., 2003; Chernozhukov et al., 2018; Zhang et al., 2019; Zhang and Bradic, 2022). More recently, Cortinovis and Caron (2025) proposed a Bayes-assisted variant of PPI. All of these contributions target fixed-width confidence intervals.
Confidence sequences were first introduced by Darling and Robbins (1967) and developed further by Robbins (1970) and Lai (1976), building on earlier work from Ville (1939) and Wald (1945). Interest has surged again in recent years (Waudby-Smith and Ramdas, 2024; Waudby-Smith et al., 2024b) motivated by applications such as A/B testing. The notion is closely linked to e-values (Ramdas et al., 2023; Ramdas and Wang, 2024). Building on the e-value framework, and on earlier work by Zrnic and Candès (2024) and Waudby-Smith and Ramdas (2024), Csillag et al. (2025) proposed an exact, time-uniform PPI method that yields confidence sequences under stronger conditions (e.g. bounded outcomes) and does not leverage prior knowledge about the prediction quality of the ML model. Furthermore, the application of their method requires an active learning setup, where, for each $t$, the observation $Z_{t}$ may be labelled with strictly positive probability. In particular, it is not applicable to deterministic sequences of observations, such as the ones describing a large initial pool of unlabelled data followed by a stream of labelled data, which is the main focus of our experiments.
In the setting of double machine learning and semiparametric inference, Dalal et al. (2024) and Waudby-Smith et al. (2024b) derive asymptotic confidence sequences for target parameters in the presence of high-dimensional nuisance components.
@@ 3 Asymptotic (Bayes-assisted) confidence sequences
In this section, we begin with some background on (asymptotic) confidence sequences (CS). Then, we show how prior information may be incorporated into asymptotic CS procedures, leading to the definition of asymptotic Bayes-assisted confidence sequences.
@@ 3.1 Background
We start by defining an exact confidence sequence (Darling and Robbins, 1967), a time-uniform analogue of classical CIs.
Definition 3.1.
(Confidence Sequence) Let $(\mathcal{C}_{\alpha,t})_{t\geq 1}$ be a sequence of random subsets of $\mathbb{R}$. For $\alpha\in(0,1)$, $(\mathcal{C}_{\alpha,t})_{t\geq 1}$ is a $1-\alpha$ confidence sequence for a fixed parameter $\mu\in\mathbb{R}$ if
$\Pr(\mu\in\mathcal{C}_{\alpha,t}\text{ \leavevmode\nobreak\ for all }t\geq 1)\geq 1-\alpha.$
We now introduce the notion of an asymptotic confidence sequence (AsympCS ) (Waudby-Smith et al., 2024b, Definition 2.1).
Definition 3.2.
(Asymptotic Confidence Sequence) Let $\alpha\in(0,1)$. Let $(\widehat{\mu}_{t})_{t\geq 1}$ be a consistent sequence of estimators of $\mu$. The sequence of random intervals $(\mathcal{C}_{\alpha,t})_{t\geq 1}$, with $\mathcal{C}_{\alpha,t}=[\widehat{\mu}_{t}-L_{t},\widehat{\mu}_{t}+U_{t}]$ and $L_{t}>0$, $U_{t}>0$, is said to be an asymptotic confidence sequence, if there exists a (usually unknown) confidence sequence $(\mathcal{C}^{\star}_{\alpha,t})_{t\geq 1}$, with $\mathcal{C}^{\star}_{\alpha,t}=[\widehat{\mu}_{t}-L_{t}^{\star},\widehat{\mu}_{t}+U_{t}^{\star}]$, such that
@@
and, almost surely as $t\to\infty$, $L_{t}/L^{\star}_{t}\to 1$ and $U_{t}/U^{\star}_{t}\to 1$.
Thus, an asymptotic CS may be regarded as an approximation of an exact CS, which is arbitrarily accurate in the limit. It is worth noting that, while classical fixed-sample asymptotic CIs rely on convergence in distribution of the scaled estimators, asymptotic confidence sequences rely on the almost sure convergence of the centered lower and upper bounds relative to those of an underlying exact CS. The following is an example of an asymptotic CS, that applies to i.i.d. data.
Theorem 3.3 (Theorem 2.2 in Waudby-Smith et al. (2024b)).
Let $(Y_{t})_{t\geq 1}$ be a sequence of i.i.d. random variables with mean $\mu$ and finite variance. For any $t\geq 1$, let $\overline{Y}_{t}$ be the sample mean, and $\widehat{\sigma}_{t}^{2}$ be the sample variance based on the first $t$ observations. For any parameter $\rho>0$, the sequence of intervals defined as
$\mathcal{C}^{\mathrm{\mathtt{N}A}}_{\alpha,t}(\overline{Y}_{t},\widehat{\sigma}_{t};\rho):=\left[\overline{Y}_{t}\pm\frac{\widehat{\sigma}_{t}}{\sqrt{t}}\sqrt{\left(1+\frac{1}{t\rho^{2}}\right)\log\left(\frac{t\rho^{2}+1}{\alpha^{2}}\right)}\right]$ (5)
forms a $(1-\alpha)-$AsympCS for $\mu$.
For the sequel, it is useful to highlight some aspects of the proof of (Waudby-Smith et al., 2024b, Theorem 2.2). First, if the random variables $(Y_{t})_{t\geq 1}$ were Gaussian with variance $\sigma^{2}$, then $\mathcal{C}^{\mathrm{\mathtt{N}A}}_{\alpha,t}(\overline{Y}_{t},\sigma;\rho)$ would be an exact CS. This follows from combining the method of mixtures for non-negative martingales with Ville’s inequality Ville (1939); Robbins (1970); Lai (1976); Howard et al. (2021). Second, the proof relies on Strassen’s strong coupling (Strassen, 1964): there exists i.i.d. Gaussian random variables $(W_{t})_{t\geq 1}$ with mean $\mu$ and variance $\mathrm{var}(Y)$ such that
$\frac{1}{t}\sum_{i=1}^{t}Y_{i}=\frac{1}{t}\sum_{i=1}^{t}W_{i}+o\left(\sqrt{\frac{\log\log t}{t}}\right)\text{ a.s. as }t\to\infty.$
Such a coupling plays a central role in constructing asymptotic confidence sequences, serving as a substitute for the CLT assumption underlying classical fixed-sample CIs. As noted in (Waudby-Smith et al., 2024b, §3.5), the construction in Theorem 3.3 extends beyond the i.i.d. case, provided a similar coupling exists.
Theorem 3.4.
@@
$\widehat{\mu}_{t}=\frac{1}{t}\sum_{i=1}^{t}W_{i}+o\left(\sqrt{\frac{\log t}{t}}\right)\text{ a.s. as }t\to\infty.$ (6)
Let $(\widehat{\sigma}^{2}_{t})_{t\geq 1}$ be a consistent sequence of estimators of $\sigma^{2}$. Then, for any parameter $\rho>0$, the sequence of intervals $(\mathcal{C}^{\mathrm{\mathtt{N}A}}_{\alpha,t}(\widehat{\mu}_{t},\widehat{\sigma}_{t};\rho))_{t\geq 1}$ forms a $(1-\alpha)$–AsympCS for $\mu$.
The asymptotic CS (5) includes a tuning parameter $\rho$, which can be chosen so as to minimise the width of the interval at a specified time $t$; see (Waudby-Smith et al., 2024b, Appendix B.2). However, this method does not allow the incorporation of prior information about the parameter of interest to yield tighter intervals when the data align with those assumptions: the width of Equation 5 is indeed independent of $\overline{Y}_{t}$.

### v2

# Prediction-Powered Inference

@@ Abstract
Given a large pool of unlabelled data and a smaller amount of labels, prediction-powered inference (PPI) leverages machine learning predictions to increase the statistical efficiency of confidence interval procedures based solely on labelled data, while preserving fixed-time validity. In this paper, we extend the PPI framework to the sequential setting, where labelled and unlabelled datasets grow over time. Exploiting Ville’s inequality and the method of mixtures, we propose prediction-powered confidence sequence procedures that are asymptotically valid uniformly over time and naturally accommodate prior knowledge on the quality of the predictions to further boost efficiency. We carefully illustrate the design choices behind our method and demonstrate its effectiveness in real and synthetic examples.
@@ 1 Introduction
Increasing the sample size of an experiment is arguably the single simplest way to improve the precision of the statistical conclusions drawn from it. However, in many fields – such as healthcare, finance, and social sciences – obtaining labelled data is often costly and time-consuming. In these settings, using machine learning (ML) models to impute additional labels represents a tempting alternative to expensive data collection, albeit at the risk of introducing bias. Prediction-powered inference (PPI) [1] is a recently introduced framework for valid statistical inference in the presence of a small labelled dataset and a large number of unlabelled examples paired with predictions from a black-box model.
@@
$\theta^{\star}=\underset{\theta\in\mathbb{R}}{\arg\min}~~\mathbb{E}[\ell_{\theta}(X,Y)],$ (1)
where $\ell_{\theta}(x,y)$ is a convex loss function parameterised by $\theta\in\mathbb{R}$. As an example, the mean $\theta^{\star}=\mathbb{E}[Y]$ is the estimand induced by the squared loss $\ell_{\theta}(x,y)=(\theta-y)^{2}/2$. For $t=1,2,\ldots$, we observe a sequence of independent random variables $Z_{t}$, either drawn from $\mathbb{P}$ (labelled sample) or from $\mathbb{P}_{X}$ (unlabelled sample), and we are provided with a black-box prediction rule $f$ that maps any input $x$ to a prediction $f(x)$.
Let $(X_{i},Y_{i})_{i\geq 1}$ and $(\widetilde{X}_{j})_{j\geq 1}$ denote the subsequence of labelled and unlabelled samples, respectively. For $n=1,2,\ldots$, let $N_{n}$ denote the number of unlabelled samples observed before the $n$th labelled one, and assume that $N_{n}\geq n$, with $N_{n}\gg n$ in typical settings. PPI constructs an (asymptotic) $1-\alpha$ confidence interval (CI) $\mathcal{C}_{\alpha,n}^{\mathrm{pp}}$ for $\theta^{\star}$, that exploits the auxiliary information encoded in $f$. To this end, under mild assumptions, $\theta^{\star}$ can be expressed as the solution to
@@
where $\ell_{\theta}^{\prime}$ is a subgradient of $\ell_{\theta}$ with respect to $\theta$. The quantity $g_{\theta}$ in Equation˜2 can be decomposed as $g_{\theta}=m_{\theta}+\Delta_{\theta}$, where
@@
where $m_{\theta}$ represents a measure of fit of the predictor, while $\Delta_{\theta}$, the rectifier, accounts for the discrepancy between the predicted outputs $f(X)$ and the true labels $Y$. If $\mathcal{C}^{g}_{\alpha,\theta,n}$ is a $(1-\alpha)$ confidence interval for $g_{\theta}$, then the PPI confidence interval $\mathcal{C}_{\alpha,n}^{\mathrm{pp}}$, defined as
@@
also achieves the desired coverage, i.e. $\Pr(\theta^{\star}\in\mathcal{C}_{\alpha,n}^{\mathrm{pp}})\geq 1-\alpha$. Constructing $\mathcal{C}^{g}_{\alpha,\theta,n}$ naturally relies on estimating $g_{\theta}$, for which PPI defines an estimator that leverages both the unlabelled data and the prediction rule $f$. The resulting method outperforms standard CI procedures based on the labelled data alone when $f$ is sufficiently accurate and $N_{n}\gg n$. Intuitively, this is because, in this case, $\Delta_{\theta}$ is close to zero, while $m_{\theta}$ can be estimated with low variance from the unlabelled data.
Crucially, coverage of the PPI CI (4) is guaranteed only at a fixed time, i.e., for a labelled sample size $n$ fixed in advance. This is undesirable in many practical settings – such as online learning, real-time monitoring, or sequential decision-making – where it is essential to continuously draw conclusions as new data arrive. In this work, we address this by proposing an anytime-valid extension of the PPI CI (4). That is, we define a confidence sequence $(\mathcal{C}_{\alpha,n}^{\mathrm{avpp}})_{n\geq 1}$, satisfying asymptotically the stronger coverage guarantee
@@
while still taking advantage of the prediction rule $f$. Analogously to standard PPI, we construct a confidence sequence $(\mathcal{C}^{g}_{\alpha,\theta,n})_{n\geq 1}$ for $g_{\theta}$ and define $\mathcal{C}_{\alpha,n}^{\mathrm{avpp}}$ through Equation˜4 for $n\geq 1$. While our approach is agnostic to the specific form of the confidence sequence $(\mathcal{C}^{g}_{\alpha,\theta,n})_{n\geq 1}$, we mainly focus on asymptotic confidence sequences [15], as they provide a versatile time-uniform analogue of standard CLT-based CIs that applies to the PPI framework above in full generality. Moreover, being based on the method of mixtures [9, 4, 8], they can readily accommodate prior information on the quality of the prediction model $f$. In particular, by means of a zero-centred prior on the rectifier $\Delta_{\theta}$, we obtain tighter confidence sequences when the predictions are good, extending the fixed-time Bayes-assisted approach of Cortinovis and Caron [6].
The remainder of the paper is organised as follows. Section˜2 reviews related work. Section˜3 provides background on (asymptotic) confidence sequences and discusses how prior information may be incorporated into their construction. Section˜4 presents PPI in the context of control-variate estimators, whose asymptotic properties are crucial for our approach to anytime-valid, Bayes-assisted PPI, which is described in Section˜5. Section˜6 demonstrates the benefits of our method on synthetic and real data. Finally, Section˜7 discusses limitations and further extensions of our approach. Proofs and additional experiments are provided in the Supplementary Material.
@@ 2 Related Work
PPI was introduced by Angelopoulos et al. [1] as a general framework for valid statistical inference with black-box machine learning predictors, and was later extended in Angelopoulos et al. [7]. Closely related ideas appear in the literatures on semi-supervised inference, missing-data methods, survey sampling, and double machine learning [8, 9, 10, 11, 12]. More recently, Cortinovis and Caron [6] proposed a Bayes-assisted variant of PPI. All of these contributions target fixed-time confidence intervals.
Confidence sequences were first introduced by Darling and Robbins [13] and developed further by Robbins and Siegmund [4] and Lai [8], building on earlier work by Ville [9] and Wald [11]. Interest has surged again in recent years [15, 15], motivated by applications such as A/B testing. The notion is closely linked to e-values [17, 17]. Building on the e-value framework, and on earlier work by Zrnic and Candès [18] and Waudby-Smith and Ramdas [15], Csillag et al. [19] proposed an exact, time-uniform PPI method that yields confidence sequences under stronger conditions (e.g., existence of bounded e-values) and does not leverage prior knowledge about the ML prediction quality. Furthermore, application of their method requires an active-learning setup in which, for each $t$, the observation $Z_{t}$ can be labelled with strictly positive probability. In particular, it is not applicable to deterministic sequences of observations, such as those describing a large initial pool of unlabelled data followed by a stream of labelled data – the main focus of our experiments.
In the setting of double machine learning and semiparametric inference, Dalal et al. [20] and Waudby-Smith et al. [15] derive asymptotic confidence sequences for target parameters in the presence of high-dimensional nuisance components.
@@ 3 Asymptotic (Bayes-assisted) confidence sequences
In this section, we begin with background on (asymptotic) confidence sequences (CS). We then show how prior information can be incorporated into asymptotic CS procedures, leading to asymptotic Bayes-assisted confidence sequences.
@@ 3.1 Background
We start by defining an exact confidence sequence [13], a time-uniform analogue of classical CIs.
Definition 1 (Confidence sequence).
Let $(\mathcal{C}_{\alpha,t})_{t\geq 1}$ be a sequence of random subsets of $\mathbb{R}$. For $\alpha\in(0,1)$, $(\mathcal{C}_{\alpha,t})_{t\geq 1}$ is a $1-\alpha$ confidence sequence for a fixed parameter $\mu\in\mathbb{R}$ if
$\Pr(\mu\in\mathcal{C}_{\alpha,t}\text{ ~for all }t\geq 1)\geq 1-\alpha.$ (5)
We now introduce the notion of an asymptotic confidence sequence (AsympCS) [15, 20].
Definition 2 (Asymptotic confidence sequence).
Let $\alpha\in(0,1)$ and $(a_{t})_{t\geq 1}$ be a real sequence such that $\lim_{t\to\infty}a_{t}=0$. Let $(\widehat{\mu}_{t})_{t\geq 1}$ be a consistent sequence of estimators of $\mu$. The sequence of random intervals $(\mathcal{C}_{\alpha,t})_{t\geq 1}$, with $\mathcal{C}_{\alpha,t}=[\widehat{\mu}_{t}-L_{t},\widehat{\mu}_{t}+U_{t}]$ and $L_{t}>0$, $U_{t}>0$, is said to be an asymptotic confidence sequence with (little-o) approximation rate $a_{t}$ if there exists a (usually unknown) confidence sequence $(\mathcal{C}^{\star}_{\alpha,t})_{t\geq 1}$, with $\mathcal{C}^{\star}_{\alpha,t}=[\widehat{\mu}_{t}-L_{t}^{\star},\widehat{\mu}_{t}+U_{t}^{\star}]$, such that
@@
and, almost surely as $t\to\infty$, $\max\{L_{t}^{\star}-L_{t},U_{t}^{\star}-U_{t}\}=o(a_{t})$.
Thus, an asymptotic CS may be regarded as an approximation of an exact CS that becomes arbitrarily accurate in the limit. It is worth noting that, while classical fixed-sample asymptotic CIs rely on convergence in distribution of the scaled estimators, asymptotic confidence sequences rely on the almost sure convergence at a given rate of the centred lower and upper bounds relative to those of an underlying exact CS. The following is an example of an asymptotic CS that applies to i.i.d. data.
Theorem 1.
Let $(Y_{t})_{t\geq 1}$ be a sequence of i.i.d. random variables with mean $\mu$ and such that $\mathbb{E}|Y_{1}|^{2+\delta}<\infty$ for some $\delta>0$. For any $t\geq 1$, let $\overline{Y}_{t}$ be the sample mean, and $\widehat{\sigma}_{t}^{2}$ be the sample variance based on the first $t$ observations. For any parameter $\rho>0$, the sequence of intervals defined as
$\mathcal{C}^{\mathrm{\tt NA}}_{\alpha,t}(\overline{Y}_{t},\widehat{\sigma}_{t};\rho):=\left[\overline{Y}_{t}\pm\frac{\widehat{\sigma}_{t}}{\sqrt{t}}\sqrt{\left(1+\frac{1}{t\rho^{2}}\right)\log\left(\frac{t\rho^{2}+1}{\alpha^{2}}\right)}\right]$ (6)
forms a $(1-\alpha)$–AsympCS with approximation rate $1/\sqrt{t\log{t}}$ for $\mu$.
For the sequel, it is useful to highlight some aspects of the proof of this theorem. First, if the random variables $(Y_{t})_{t\geq 1}$ were Gaussian with variance $\sigma^{2}$, then $\mathcal{C}^{\mathrm{\tt NA}}_{\alpha,t}(\overline{Y}_{t},\sigma;\rho)$ would be an exact CS. This follows from combining the method of mixtures for nonnegative martingales with Ville’s inequality [9, 4, 8, 14]. Second, the proof relies on KMT strong coupling [2, 3]: there exists i.i.d. Gaussian random variables $(W_{t})_{t\geq 1}$ with mean $\mu$ and variance $\mathrm{var}(Y)$ such that
$\frac{1}{t}\sum_{i=1}^{t}Y_{i}=\frac{1}{t}\sum_{i=1}^{t}W_{i}+o\left(\frac{1}{\sqrt{t\log t}}\right)\text{ a.s. as }t\to\infty.$
Such a coupling plays a central role in constructing asymptotic confidence sequences, serving as a substitute for the CLT assumption underlying classical fixed-sample CIs. The construction in Theorem˜1 extends beyond the i.i.d. case, provided a similar coupling exists.
Theorem 2.
@@
$\widehat{\mu}_{t}=\frac{1}{t}\sum_{i=1}^{t}W_{i}+o\left(\frac{1}{\sqrt{t\log t}}\right)\text{ a.s. as }t\to\infty.$ (7)
Let $(\widehat{\sigma}^{2}_{t})_{t\geq 1}$ be a consistent sequence of estimators of $\sigma^{2}$ with $|\widehat{\sigma}_{t}-\sigma|=o\left(\frac{1}{\log t}\right)$ a.s. Then, for any parameter $\rho>0$, the sequence of intervals $(\mathcal{C}^{\mathrm{\tt NA}}_{\alpha,t}(\widehat{\mu}_{t},\widehat{\sigma}_{t};\rho))_{t\geq 1}$ forms a $(1-\alpha)$–AsympCS with approximation rate $1/\sqrt{t\log{t}}$ for $\mu$.
The asymptotic CS (6) includes a tuning parameter $\rho$, which can be chosen so as to minimise the width of the interval at a specified time $t$; see [15, Appendix B.2]. However, this method does not allow the incorporation of prior information about the parameter of interest to yield tighter intervals when the data align with those assumptions: the width of Equation˜6 is indeed independent of $\overline{Y}_{t}$.